A 25.0 L tank contains 2.0 kg of neon gas with a pressure of 98.0 atm at . How many kilograms of can fill the tank under the same pressure and temperature?
1
Expert's answer
2020-12-07T07:22:56-0500

Solution.

"p \times V = n \times R \times T"

"n = \frac{m}{M}"

n = 100 mol

"T = \frac{p \times V}{n \times R}"

T = 298.73 K

"n1 = \frac{101325 \times 0.025}{8.31 \times 298.73}"

n1 = 1.02 mol

"m(Ne) = n \times M(Ne)"

m(Ne) = 0.02 kg

Answer:

m(Ne) = 0.02 kg